OpenStudy (anonymous):

in 2006, there were 12,440 starbucks stores worldwide. this is approximately 1126% more stores than 10 years earlier. how many starbucks stores were there in 1996?

hero (hero):

I'll be back later.

OpenStudy (anonymous):

Let y = number of stores in 2006; Let x = number of stores in 1996; y = (1126/100)*x Solve for x From the question given, y = 12440; 12440 = (1126/100)x Multiply both sides by 100 1244000 = 1126x Divide both sides by 1126 x = 1105 stores

hero (hero):

I was trying to do something similar to that but got the numbers mixed up

hero (hero):

is/of = %/100 12440/x = 1126/100 1244000 = 1126x 1244000/1126 = x 1104.8 = x Technically, you're supposed to round down in this case since there can't be 1104.8 stores. So 1104 stores.